SOFSEM ’94 – Milovy, The Czech Republic, 27. 11. – 9. 12. 1994

Typesetting Articles for the SOFSEM Proceedings^{∗}

Petr Sojka

Institute of Computer Science, Masaryk University Brno Burešova 20, CZ–602 00 Brno, The Czech Republic

                            Email: sojka@muni.cz.

Abstract: This article gives advice to the preparation of camera-ready contributions for the SOFSEM

Proceedings.

It describes the recommended instructions, usage and, if needed, the implementation of the sofsem LaTEX style option for LaTEX users.

Keywords: SOFSEM, TEX^{†}, LaTEX, typesetting, camera ready, instructions for authors

Motto: “How do we persuade authors that spreading fonts accross the page like peanut butter across a hot toast is not necessarily the route to typographic excellence?” Peter Flynn

1 Preface

The SOFSEM Proceedings, like the other conference proceedings, is rushed to publication as quickly as possible so that the material will reach the SOFSEM participants (and other interested readers) just in conference time. Reproducing author-prepared copy eliminates the considerable bother and expense of typesetting, proofreading and corrections. The published document should be compact, uniform in appearance, and readable.

The authors are adviced to send camera ready copy of their articles according to these instructions (style file). To achieve uniformity, the author is provided with this “model paper”. This article (which has itself been produced by the technique it describes) introduces the sofsem.sty LaTEX style option, designed to prepare Proceedings articles using LaTEX [2], well-known macro package for TEX typesetting system [1].

2 Instructions for authors

The submitted text should be prepared in the camera-ready format required for publication (printed on a laser-printer). Authors are suggested to use a word processing equipment (LaTEX) and they should observe the following instructions at the preparation of their text:

Language: The text should be in English.

†TEX is a trademark of the American Mathematical Society.

∗Instructions for authors of SOFSEM Proceedings contributions

P. Sojka: Typesetting Articles for the SOFSEM Proceedings

Sheet size: Produce your text on plain white DIN-A4 paper. The printout must be sharp, black and even (not grey).

Fonts and format: The text should be printed using 11-point proportional fonts (Adobe Times-Roman preferred) in the frame 160 mm wide and 240 mm long (no footers), preferably in two columns format.

Heading: The heading of the contribution should follow appearence of this article. The text follows immediately the heading on the first page.

Running heads: Running heads should follow appearance of this article in

the form <author>:<article title>

on the left (verso) page, and

<section number>.<section title>

on the right (recto) page.

Titles & emphasizing: Titles of the lecture, sections and subsections, should not be underlined. Follow emphasizing principles used in this article.

Page numbering: A pencil should be used for numbering the pages, on back sides. Do not number pages on front side! Leave enough space for page numbers at the page bottom.

Citations: Citations should be in the form [1], [2]. References should be written according to bibliography of this article [3, 4, 5]. For BibTEX users bibliography style sofsem.bst is provided.

Page limit: The text of your contribution should not exceed specified number of pages. Page limit is 10 pages for one 75 minute lecture (e. g. 20 pages for 2lecture talk, 30 pages for 3-lecture talk)

3 Deadline and Address

One camera-ready copy of invited papers should reach the SOFSEM secretary not later then by 20^{th} September.^{1} We prefer to receive both the camera-ready paper and (if possible) the postscript file at the address:

SOFSEM ’94

Institute of Computer Science MU Burešova 20 CZ–602 00 Brno Czech Republic

                                         Email: sofsem@muni.cz

4 Support for TEX users

Typesetting using this style file^{2} is actively supported by organizers of the SOFSEM to achieve uniformity of the SOFSEM Proceedings. Latest version of the macros is available

                                    by ftp from ftp.muni.cz (login anonymous,

password your email address), in the directory

                                    pub/tex/styles/packages/sofsem.

Any questions regarding submissions or about SOFSEM in general should be sent to sofsem@muni.cz. Any questions, suggestions, complaints, bugs regarding this style should be sent to author or to the address

                                    TeX-support@muni.cz.

References

[1] Knuth, D. E.: The TEXBook. Addison-Wesley, Reading, Massachusets, 1984, 484 p.

[2] Lamport, L.: LaTEX: Document preparation System. Addison-Wesley, 1986, 242 p.

[3] Rabin, M.O. – Scott, D.: Finite automata and their decision problems. IBM Journal of Research and Development, 5, 1976, No. 2, pp 114–125

[4] Eilenberg, S.: Automata, Languages, and Machines. Vol. A, Academic Press, New York 1974

[5] Clarke, L.: A system to generate test data. Tech. Rep. University of Colorado, Boulder, Februar 1975

1The sooner you send your contribution the better!

2This file is both LaTEX style file and instructions for authors. File has version number v94.0, was last revised 94/06/03, documentation is dated 94/06/03.

5. Running head example

5 Running head example

This page is included just to show running heads on recto (odd-numbered) page.